OpenStudy (solomonzelman):

Can you proof this identity briefly enough? without posting links please.

OpenStudy (solomonzelman):

\[\huge\color{blue}{\huge {\bbox[5pt,lime,border:2px solid purple]{Log_27=\frac{Log_{10}7}{Log_{10}2}}}}\]

OpenStudy (solomonzelman):

I am leaving for an hour..

OpenStudy (anonymous):

change of base theorem I suck at proofs :(

OpenStudy (mathstudent55):

\( \log_2 7 = x \iff 2^x = 7\) \(\log_{10} 7 = y \iff 10^y = 7\) \(y = \log_{10} 7 \) \(~~= \log_{10}{2^x} \) \(~~= x\log_{10}{2}\) \( y= \log_2 7 \log_{10} 2\) \( \log_{10} 7= \log_2 7 \log_{10} 2\) \( \dfrac{\log_{10}7 }{\log_{10} 2} = \log_2 7\)

OpenStudy (solomonzelman):

WOW! NICE AND CLEAR!

OpenStudy (mathstudent55):

wlcm

OpenStudy (solomonzelman):

@jim_thompson5910 @phi @myininaya can you give math student a medal for appreciation?

OpenStudy (solomonzelman):

Really awesome, and brief I like that, and without unnecessary words...

OpenStudy (anonymous):

how do you give medals?

OpenStudy (mathstudent55):

It's just the definition of logs, a property of logs, and substitution. Then a little algebra to finish it off.

myininaya (myininaya):

Yes. You can give medals to students. I do sometimes when they put forth a good effort.